OpenStudy (anonymous):

exact value for cos^-1(cos5pi/4)

hartnn (hartnn):

hint : cos^-1 cos (theta) = theta

OpenStudy (anonymous):

\[\cos^-1 ( \cos \Pi + \Pi/4)\]

OpenStudy (anonymous):

\[\cos^-1 ( -\cos \Pi/4)\]

OpenStudy (anonymous):

= \[-\Pi/4 = -45\]

OpenStudy (lgbasallote):

you have to express co (5 pi/4) in reference angles before applying the rule cos^-1 (cos x) = x

OpenStudy (anonymous):

What is the color of Napoleon's white horse?